# Accelerated Parallel Tempering via Neural Transports

### v3

@@ B.2 Proof of Proposition˜3
Suppose that ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$ arise as $K$-step Euler-Maruyama discretizations of the following SDEs respectively
$d\accentset{\rightharpoonup}{X}^{n-1}_{t}$ $=f(t,\accentset{\rightharpoonup}{X}^{n-1}_{t})dt+\sigma dB_{t},$ (20)
$d\accentset{\leftharpoonup}{X}^{n}_{t}$ $=b(t,\accentset{\leftharpoonup}{X}^{n}_{t})dt+\sigma dB_{t},$ (21)
@@
where $r(p,q)=\|p\times q-q\times p\|_{\mathrm{TV}}$ and ${\mathbb{P}}_{\infty}^{n-1}$ and ${\mathbb{Q}}_{\infty}^{n}$ are respectively the full path measures of Equation˜20 and Equation˜21.
@@
Let ${\mathbb{P}}_{\infty|K}^{n-1}$ and ${\mathbb{Q}}_{\infty|K}^{n}$ be the restrictions of the path measures ${\mathbb{P}}_{\infty}^{n-1}$ and ${\mathbb{Q}}_{\infty}^{n}$ to the time discretization points. (As such, ${\mathbb{P}}_{\infty|K}^{n-1}$ and ${\mathbb{Q}}_{\infty|K}^{n}$ are defined on the same space as ${\mathbb{P}}_{K}^{n-1}$ and ${\mathbb{Q}}_{K}^{n}$.) We have
@@
where the first inequality is elementary and the second follows from Proposition˜6. By data processing inequality
@@
Equation˜24 and Equation˜25 establish Equation˜23. On the other hand, putting
$R_{K}(x_{k})$ $:=\frac{\mathrm{d}({\mathbb{Q}}_{K}^{n}\times{\mathbb{P}}_{K}^{n-1})}{\mathrm{d}({\mathbb{P}}_{K}^{n-1}\times{\mathbb{Q}}_{K}^{n})}(x_{k});$
$R_{\infty}(x_{\infty})$ $:=\frac{\mathrm{d}({\mathbb{Q}}_{\infty}^{n}\times{\mathbb{P}}_{\infty}^{n-1})}{\mathrm{d}({\mathbb{P}}_{\infty}^{n-1}\times{\mathbb{Q}}_{\infty}^{n})}(x_{\infty});$
$\varphi(\alpha)$ $:=1-\min(1,\alpha)$
@@
In addition let $D_{K}:\mathcal{C}[0,1]\to(\mathbb{R}^{d})^{K+1}$ be the discretisation operator which takes a continuous path $x_{\infty}$ and extract its values at $K+1$ discretisation points. Then
@@
$\begin{split}\mathbb{E}_{{\mathbb{P}}_{\infty|K}^{n-1}\times{\mathbb{Q}}_{\infty|K}^{n}}[\varphi\circ R_{K}]&=\mathbb{E}_{{\mathbb{P}}_{\infty}^{n-1}\times{\mathbb{Q}}_{\infty}^{n}}[\varphi\circ R_{K}\circ D_{K}]\\ &\overset{K\to\infty}{\longrightarrow}\mathbb{E}_{{\mathbb{P}}_{\infty}^{n-1}\times{\mathbb{Q}}_{\infty}^{n}}[\varphi\circ R_{\infty}]=r({\mathbb{P}}_{\infty}^{n-1},{\mathbb{Q}}_{\infty}^{n})\end{split}$ (28)
by dominated convergence theorem, taking into account the pathwise convergence of the Euler-Maruyama schemes (Kloeden & Neuenkirch (2007); Gyöngy (1998), see also Berner et al. (2025, Lemma B.7.a)). Equation˜26, Equation˜27, and Equation˜28 entail Equation˜22, finishing the proof.
@@
$\|{\mathbb{P}}^{*}_{K}-{{\mathbb{P}}}_{K}\|_{\mathrm{TV}}\in\mathcal{O}(\frac{1}{\sqrt{K}}).$
@@
Consider the continuous-time extension $(\hat{X}_{t})_{t\in[0,1]}$ of the $K$-step Euler-Maruyama discretization of the SDE in Equation˜29 given by
@@
$\leq 2G(t-t_{k})\int_{t_{k}}^{t}\mathbb{E}(1+\|X_{s}\|)^{2}ds+2d\sigma^{2}(t-t_{k})$
$\leq 4G(t-t_{k})\int_{t_{k}}^{t}(1+\mathbb{E}\|X_{s}\|^{2})ds+2d\sigma^{2}(t-t_{k})$
@@
$\leq\frac{1}{2\sigma^{2}}\sum_{k=0}^{K-1}\frac{8}{3}GL^{2}(1+M)(t_{k+1}-t_{k})^{3}+2d\sigma^{2}L^{2}(t_{k+1}-t_{k})^{2}+\frac{2}{3}MB^{2}(t_{k+1}-t_{k})^{3}$
$\leq\frac{4}{3\sigma^{2}}GL^{2}(1+M)\frac{1}{K^{2}}+dL^{2}\frac{1}{K}+\frac{1}{3\sigma^{2}}MB^{2}\frac{1}{K^{2}}.$
@@
B.3 Further Details on Scaling with Parallel Chains
B.3.1 Regularity Assumptions
Suppose $\pi^{\beta}(x)=\exp(-U^{\beta}(x))/Z_{\beta}$, where $Z_{\beta}=\int_{\mathcal{X}}\exp(-U^{\beta}(x))\mathrm{d}x$. For $\bm{\beta}=(\beta,\beta^{\prime})$, suppose ${\mathbb{P}}^{\bm{\beta}}_{K}(\mathrm{d}x_{0:K})$ and ${\mathbb{Q}}^{\bm{\beta}}_{K}(\mathrm{d}x_{0:K})$ are mutually absolutely continuous path measures on $\mathcal{X}^{K+1}$ with marginals ${\mathbb{P}}^{\bm{\beta}}(x_{0})=\pi^{\beta}(\mathrm{d}x_{0})$, and ${\mathbb{Q}}^{\bm{\beta}}(\mathrm{d}x_{K})=\pi^{\beta^{\prime}}(\mathrm{d}x_{K})$ with work functional $W^{\bm{\beta}}:\mathcal{X}^{K+1}\mapsto\mathbb{R}$,
$W^{\bm{\beta}}(x_{0:K})=\Delta F_{\bm{\beta}}-\log\frac{\mathrm{d}{\mathbb{P}}^{\bm{\beta}}}{\mathrm{d}{\mathbb{Q}}^{\bm{\beta}}}(x_{0:K}),$
with $\Delta F_{\bm{\beta}}:=\log Z_{\beta^{\prime}}-\log Z_{\beta}$.
@@ B.3.1 Regularity Assumptions
  • $\bm{\beta}\mapsto g^{\bm{\beta}}(x_{0:K},x_{0:K})$ is continuously differentiable,
@@ B.3.2 Proof of Theorem˜2
$\alpha^{\bm{\beta}}(x_{0:K},x_{0:K}^{\prime})$ $=\exp\left(\min\left\{0,\Delta W^{\bm{\beta}}_{K}(x_{0:K},x_{0:K}^{\prime})\right\}\right)$
and $\Delta W^{\bm{\beta}}:\mathcal{X}^{K+1}\times\mathcal{X}^{K+1}\to\mathbb{R}$ is the change in work,
$\Delta W^{\bm{\beta}}(x_{0:K},x_{0:K}^{\prime}):=W^{\bm{\beta}}(x_{0:K}^{\prime})-W^{\bm{\beta}}(x_{0:K}).$
@@
Given an annealing schedule $0=\beta_{0}<\cdots<\beta_{N}=1$, let $\bm{\beta}_{n}=(\beta_{n-1},\beta_{n})$ and $\Delta\beta_{n}=\beta_{n}-\beta_{n-1}$. It follows from Lemma˜3 that the sum of the rejection satisfies,
@@
We will first approximate the first term in Equation˜30. Note that since $\Delta W^{\beta}=0$ and $\partial_{\beta^{\prime}}\Delta W^{\bm{\beta}}=\lim_{\beta^{\prime}\to\beta}\Delta W^{\bm{\beta}}/\Delta\beta=:\Delta\dot{W}^{\bm{\beta}}$, we have that $\dot{\alpha}^{\beta}$ satisfies:
@@
For the second term in Equation˜30, we have
@@
By Lemma˜2, $\beta\mapsto\lambda_{\beta}$ is continuously differentiable. Since $\lambda_{\beta}\Delta\beta$ is a right Riemann sum for the integral of $\lambda_{\beta}$ with error:
@@
Appendix C Further Details on Design-Space for Accelerated PT
C.1 Further Details on Flow APT
C.1.1 Work Formula for Deterministic flows

### v4

@@ B.3 Proof of Proposition˜3
Consider two SDEs
$\accentset{\rightharpoonup}{X}^{n-1}(0)\sim\pi^{n-1},$ $\quad d\accentset{\rightharpoonup}{X}^{n-1}(u)=f(u,\accentset{\rightharpoonup}{X}^{n-1}(u))du+\sigma dB_{u}\text{ for }u\in[0,1]$ (15)
$\accentset{\leftharpoonup}{X}^{n}(1)\sim\pi^{n},$ $\quad d\accentset{\leftharpoonup}{X}^{n}(u)=b(u,\accentset{\leftharpoonup}{X}^{n}(u))dt+\sigma dB^{\prime}_{u}\text{ for }u\in[0,1];$ (16)
@@
where $r(p,q)=\|p\times q-q\times p\|_{\mathrm{TV}}$.
@@
Let ${\mathbb{P}}_{\infty|K}^{n-1}$ and ${\mathbb{Q}}_{\infty|K}^{n}$ be the restrictions of the path measures ${\mathbb{P}}_{\infty}^{n-1}$ and ${\mathbb{Q}}_{\infty}^{n}$ to the time discretisation points. (As such, ${\mathbb{P}}_{\infty|K}^{n-1}$ and ${\mathbb{Q}}_{\infty|K}^{n}$ are defined on the same space as ${\mathbb{P}}_{K}^{n-1}$ and ${\mathbb{Q}}_{K}^{n}$.) We have
@@
where the first inequality is elementary and the second follows from Proposition˜4 below. By data processing inequality
@@
Equation˜19 and Equation˜20 establish Equation˜18. On the other hand, putting
$R_{K}(x_{k},y_{k})$ $:=\frac{\mathrm{d}({\mathbb{Q}}_{K}^{n}\times{\mathbb{P}}_{K}^{n-1})}{\mathrm{d}({\mathbb{P}}_{K}^{n-1}\times{\mathbb{Q}}_{K}^{n})}(x_{k},y_{k});$ (21)
$R_{\infty}(x_{\infty},y_{\infty})$ $:=\frac{\mathrm{d}({\mathbb{Q}}_{\infty}^{n}\times{\mathbb{P}}_{\infty}^{n-1})}{\mathrm{d}({\mathbb{P}}_{\infty}^{n-1}\times{\mathbb{Q}}_{\infty}^{n})}(x_{\infty},y_{\infty});$ (22)
$\varphi(\alpha)$ $:=1-\min(1,\alpha)$ (23)
@@
In addition define the discretisation operator
$D_{K}:\mathcal{C}[0,1]$ $\to(\mathbb{R}^{d})^{K+1};$
$x_{\infty}$ $\mapsto x_{\infty|K}=[x_{\infty}(0),x_{\infty}(1/K),\ldots,x_{\infty}(1)]$ (26)
@@
which takes a continuous path $x_{\infty}$ and extract its values at $K+1$ discretisation points. Let
$\tilde{D}_{K}(x_{\infty},y_{\infty})=(D_{K}(x_{\infty}),D_{K}(y_{\infty})).$ (27)
Then
@@
$\begin{split}\mathbb{E}_{{\mathbb{P}}_{\infty|K}^{n-1}\times{\mathbb{Q}}_{\infty|K}^{n}}[\varphi\circ R_{K}]&=\mathbb{E}_{{\mathbb{P}}_{\infty}^{n-1}\times{\mathbb{Q}}_{\infty}^{n}}[\varphi\circ R_{K}\circ\tilde{D}_{K}]\\ &\overset{K\to\infty}{\longrightarrow}\mathbb{E}_{{\mathbb{P}}_{\infty}^{n-1}\times{\mathbb{Q}}_{\infty}^{n}}[\varphi\circ R_{\infty}]=r({\mathbb{P}}_{\infty}^{n-1},{\mathbb{Q}}_{\infty}^{n})\end{split}$ (28)
by Lemma˜2 below. Finally, Equation˜24, Equation˜25, and Equation˜28 entail Equation˜17, finishing the proof. ∎
Lemma 2.
We have
$\mathbb{E}_{{\mathbb{P}}_{\infty}^{n-1}\times{\mathbb{Q}}_{\infty}^{n}}[\varphi\circ R_{K}\circ\tilde{D}_{K}]\overset{K\to\infty}{\longrightarrow}\mathbb{E}_{{\mathbb{P}}_{\infty}^{n-1}\times{\mathbb{Q}}_{\infty}^{n}}[\varphi\circ R_{\infty}]$
where the quantities $R_{K}$ and $R_{\infty}$ are defined in Equation˜21 and Equation˜22; the function $\varphi$ is given by Equation˜23; and the operators $D_{K}$ and $\tilde{D}_{K}$ are defined in Equation˜26 and Equation˜27.
Proof.
Introduce the reference measure $\mathbb{M}(\mathrm{d}x_{\infty})$ as the solution of the stationary SDE
$\mathrm{d}X_{u}=-\frac{\sigma^{2}}{2}X_{u}\mathrm{d}u+\sigma\mathrm{d}B_{u},\quad u\in[0,1]\text{ and }X_{0}\sim\mathcal{N}(0,\operatorname{Id}).$
Let $\mathbb{M}^{f}_{K}(\mathrm{d}z)$ be the discrete measure defined on $(\mathbb{R}^{d})^{K+1}$ and given by
$z_{0}\sim\mathcal{N}(0,\operatorname{Id}),\quad z_{k+1}|z_{k}\sim\mathcal{N}(z_{k}-\frac{\sigma^{2}}{2K}z_{k};\frac{\sigma^{2}}{K}\operatorname{Id}).$
Similarly let $\mathbb{M}^{b}_{K}(\mathrm{d}z)$ be the discrete measure defined on $(\mathbb{R}^{d})^{K+1}$ and given by
$z_{K}\sim\mathcal{N}(0,\operatorname{Id}),\quad z_{k}|z_{k+1}\sim\mathcal{N}(z_{k+1}-\frac{\sigma^{2}}{2K}z_{k+1};\frac{\sigma^{2}}{K}\operatorname{Id}).$
We first verify the three statements:
$\frac{\mathrm{d}{\mathbb{Q}}_{K}^{n}}{\mathrm{d}\mathbb{M}^{b}_{K}}(D_{K}x_{\infty})$ $\overset{\mathbb{P}}{\longrightarrow}\frac{\mathrm{d}{\mathbb{Q}}_{\infty}^{n}}{\mathrm{d}\mathbb{M}}(x_{\infty})\text{ under }{\mathbb{Q}}_{\infty}^{n};$ (29)
$\frac{\mathrm{d}{\mathbb{P}}_{K}^{n-1}}{\mathrm{d}\mathbb{M}^{f}_{K}}(D_{K}x_{\infty})$ $\overset{\mathbb{P}}{\longrightarrow}\frac{\mathrm{d}{\mathbb{P}}_{\infty}^{n-1}}{\mathrm{d}\mathbb{M}}(x_{\infty})\text{ under }{\mathbb{P}}_{\infty}^{n-1};$ (30)
$\frac{\mathrm{d}\mathbb{M}^{f}_{K}}{\mathrm{d}\mathbb{M}^{b}_{K}}(D_{K}x_{\infty})$ $\overset{\mathbb{P}}{\longrightarrow}1\text{ under either }{\mathbb{Q}}_{\infty}^{n}\text{ or }{\mathbb{P}}_{\infty}^{n-1}.$ (31)
The stationary measure $\mathbb{M}$ admits a tractable forward and backward form. The Girsanov theorem can therefore be applied to derive a closed-form formula for $\mathrm{d}{\mathbb{Q}}_{\infty}^{n}/\mathrm{d}\mathbb{M}$ and $\mathrm{d}{\mathbb{P}}_{\infty}^{n-1}/\mathrm{d}\mathbb{M}$. The three statements are then straightforwardly verified via the convergence in probability of the Euler approximation of an Itô (forward or backward) stochastic integral (Le Gall, 2016, Proposition 5.9). For example, for Equation˜30 we have
$\frac{\mathrm{d}{\mathbb{P}}_{K}^{n-1}}{\mathrm{d}\mathbb{M}^{f}_{K}}(D_{K}x_{\infty})$ $=\frac{\pi^{n-1}(x_{\infty}(0))\prod_{k=0}^{K-1}\mathcal{N}(x_{\infty}(\frac{k+1}{K})|x_{\infty}(\frac{k}{K})+\frac{1}{K}f(\frac{k}{K},x_{\infty}(\frac{k}{K}));\frac{\sigma^{2}}{K}\operatorname{Id})}{\mathcal{N}(x_{\infty}(0)|0,\operatorname{Id})\prod_{k=0}^{K-1}\mathcal{N}(x_{\infty}(\frac{k+1}{K})|x_{\infty}(\frac{k}{K})-\frac{\sigma^{2}}{2K}x_{\infty}(\frac{k}{K});\frac{\sigma^{2}}{K}\operatorname{Id})}$
$=\frac{\pi^{n-1}(x_{\infty}(0))}{\mathcal{N}(x_{\infty}(0)|0,\operatorname{Id})}\exp\left\{\sum_{k=0}^{K-1}\left\langle\frac{f(\frac{k}{K},x_{\infty}(\frac{k}{K}))}{\sigma^{2}}+\frac{x_{\infty}(\frac{k}{K})}{2},x_{\infty}(\frac{k+1}{K})-x_{\infty}(\frac{k}{K})\right\rangle\right\}\times$
[…]
$\overset{\mathbb{P}}{\longrightarrow}\frac{\pi^{n-1}(x_{\infty}(0))}{\mathcal{N}(x_{\infty}(0)|0,\operatorname{Id})}\exp\left\{\int_{0}^{1}\left\langle\frac{f(u,x_{\infty}(u))}{\sigma^{2}}+\frac{x_{\infty}(u)}{2},\mathrm{d}x_{\infty}(u)\right\rangle\mathrm{d}u\right\}\times$
[…]
$=\frac{\mathrm{d}{\mathbb{P}}_{\infty}^{n-1}}{\mathrm{d}\mathbb{M}}(x_{\infty}).$
Now recall that any sequence of variables converging in probability admits a subsequence converging almost-surely. (The measures ${\mathbb{P}}_{\infty}^{n-1}$ and ${\mathbb{Q}}_{\infty}^{n}$ being mutually absolutely continuous, almost sure convergence can be understood w.r.t. any of them.) Given any sequence of increasing values of $K$, applying that remark sequentially three times through Equation˜29, Equation˜30 and Equation˜31 produces a subsequence $(K_{\ell})$ such that
$\frac{\mathrm{d}{\mathbb{Q}}_{K_{\ell}}^{n}}{\mathrm{d}\mathbb{M}^{b}_{K_{\ell}}}(D_{K_{\ell}}x_{\infty})$ $\overset{\text{a.s.}}{\longrightarrow}\frac{\mathrm{d}{\mathbb{Q}}_{\infty}^{n}}{\mathrm{d}\mathbb{M}}(x_{\infty});$
$\frac{\mathrm{d}{\mathbb{P}}_{K_{\ell}}^{n-1}}{\mathrm{d}\mathbb{M}^{f}_{K_{\ell}}}(D_{K_{\ell}}x_{\infty})$ $\overset{\text{a.s.}}{\longrightarrow}\frac{\mathrm{d}{\mathbb{P}}_{\infty}^{n-1}}{\mathrm{d}\mathbb{M}}(x_{\infty});$
$\frac{\mathrm{d}\mathbb{M}^{f}_{K_{\ell}}}{\mathrm{d}\mathbb{M}^{b}_{K_{\ell}}}(D_{K_{\ell}}x_{\infty})$ $\overset{\text{a.s.}}{\longrightarrow}1.$
Combining the three equations above
$\frac{\mathrm{d}{\mathbb{Q}}^{n}_{K_{\ell}}}{\mathrm{d}{\mathbb{P}}^{n-1}_{K_{\ell}}}(D_{K_{\ell}}x_{\infty})\overset{\text{a.s.}}{\longrightarrow}\frac{\mathrm{d}{\mathbb{Q}}_{\infty}^{n}}{\mathrm{d}{\mathbb{P}}_{\infty}^{n-1}}(x_{\infty})$ (32)
and, by symmetry and by the subsequence extraction trick again, we can consider w.l.o.g. that
$\frac{\mathrm{d}{\mathbb{P}}^{n-1}_{K_{\ell}}}{\mathrm{d}{\mathbb{Q}}^{n}_{K_{\ell}}}(D_{K_{\ell}}y_{\infty})\overset{\text{a.s.}}{\longrightarrow}\frac{\mathrm{d}{\mathbb{P}}_{\infty}^{n-1}}{\mathrm{d}{\mathbb{Q}}_{\infty}^{n}}(y_{\infty}).$ (33)
Since
$R_{K}\circ\tilde{D}_{K}(x_{\infty},y_{\infty})$ $=\frac{\mathrm{d}{\mathbb{Q}}_{K}^{n}}{\mathrm{d}{\mathbb{P}}_{K}^{n-1}}(D_{K}x_{\infty})\frac{\mathrm{d}{\mathbb{P}}_{K}^{n-1}}{\mathrm{d}{\mathbb{Q}}_{K}^{n}}(D_{K}y_{\infty}),$
$R_{\infty}(x_{\infty},y_{\infty})$ $=\frac{\mathrm{d}{\mathbb{Q}}_{\infty}^{n}}{\mathrm{d}{\mathbb{P}}_{\infty}^{n-1}}(x_{\infty})\frac{\mathrm{d}{\mathbb{P}}_{\infty}^{n-1}}{\mathrm{d}{\mathbb{Q}}_{\infty}^{n}}(y_{\infty}),$
the dominated convergence theorem and the boundedness of $\varphi$ gives
$\mathbb{E}_{{\mathbb{P}}_{\infty}^{n-1}\times{\mathbb{Q}}_{\infty}^{n}}[\varphi\circ R_{K_{\ell}}\circ\tilde{D}_{K_{\ell}}]\overset{\ell\to\infty}{\longrightarrow}\mathbb{E}_{{\mathbb{P}}_{\infty}^{n-1}\times{\mathbb{Q}}_{\infty}^{n}}[\varphi\circ R_{\infty}].$
Since we are always able to extract such a subsequence $(K_{\ell})$ given any increasing sequence $K$, the lemma is proved. ∎
@@
$\|{\mathbb{P}}^{*}_{K}-{{\mathbb{P}}}_{K}\|_{\mathrm{TV}}\in\mathcal{O}\left(\frac{1}{\sqrt{K}}\right).$
@@
Consider the continuous-time extension $(\hat{X}_{t})_{t\in[0,1]}$ of the $K$-step Euler-Maruyama discretization of the SDE in Equation˜34 given by
@@
$\leq 2G^{2}(t-t_{k})\int_{t_{k}}^{t}\mathbb{E}(1+\|X_{s}\|)^{2}ds+2d\sigma^{2}(t-t_{k})$
$\leq 4G^{2}(t-t_{k})\int_{t_{k}}^{t}(1+\mathbb{E}\|X_{s}\|^{2})ds+2d\sigma^{2}(t-t_{k})$
@@
$\leq\frac{1}{2\sigma^{2}}\sum_{k=0}^{K-1}\frac{8}{3}G^{2}L^{2}(1+M)(t_{k+1}-t_{k})^{3}+2d\sigma^{2}L^{2}(t_{k+1}-t_{k})^{2}+\frac{2}{3}B^{2}(t_{k+1}-t_{k})^{3}$
$\leq\frac{4}{3\sigma^{2}}G^{2}L^{2}(1+M)\frac{1}{K^{2}}+dL^{2}\frac{1}{K}+\frac{1}{3\sigma^{2}}B^{2}\frac{1}{K^{2}}.$
@@
B.4 Further Details on Scaling with Parallel Chains
B.4.1 Regularity Assumptions
Suppose $\pi^{\beta}(x)=\exp(-U^{\beta}(x))/Z_{\beta}$, where $Z_{\beta}=\int_{\mathcal{X}}\exp(-U^{\beta}(x))\mathrm{d}x$. For $\bm{\beta}=(\beta,\beta^{\prime})$, suppose ${\mathbb{P}}^{\bm{\beta}}_{K}(\mathrm{d}x_{0:K})$ and ${\mathbb{Q}}^{\bm{\beta}}_{K}(\mathrm{d}x_{0:K})$ are mutually absolutely continuous path measures on $\mathcal{X}^{K+1}$ with marginals ${\mathbb{P}}^{\bm{\beta}}(x_{0})=\pi^{\beta}(\mathrm{d}x_{0})$, and ${\mathbb{Q}}^{\bm{\beta}}(\mathrm{d}x_{K})=\pi^{\beta^{\prime}}(\mathrm{d}x_{K})$ with weight functional $w^{\bm{\beta}}:\mathcal{X}^{K+1}\mapsto\mathbb{R}$,
$w^{\bm{\beta}}(x_{0:K})=\frac{Z_{\beta^{\prime}}}{Z_{\beta}}\frac{\mathrm{d}{\mathbb{Q}}^{\bm{\beta}}}{\mathrm{d}{\mathbb{P}}^{\bm{\beta}}}(x_{0:K}),$
@@ B.4.1 Regularity Assumptions
  • $\bm{\beta}\mapsto g^{\bm{\beta}}(x_{0:K},x^{\prime}_{0:K})$ is continuously differentiable,
@@ B.4.2 Proof of Theorem˜2
$\alpha^{\bm{\beta}}(x_{0:K},x_{0:K}^{\prime})$ $=\min\left\{1,\frac{w^{\bm{\beta}}_{K}(x_{0:K})}{w^{\bm{\beta}}_{K}(x_{0:K}^{\prime})}\right\}=\exp\left(\min\left\{0,\Delta W^{\bm{\beta}}_{K}(x_{0:K},x_{0:K}^{\prime})\right\}\right)$
where $\Delta W^{\bm{\beta}}:\mathcal{X}^{K+1}\times\mathcal{X}^{K+1}\to\mathbb{R}$ is the change in log-weight $W^{\bm{\beta}}(x_{0:K}):=\log w^{\bm{\beta}}(x_{0:K})$,
$\Delta W^{\bm{\beta}}(x_{0:K},x_{0:K}^{\prime}):=W^{\bm{\beta}}(x_{0:K})-W^{\bm{\beta}}(x_{0:K}^{\prime})$
@@
Given an annealing schedule $0=\beta_{0}<\cdots<\beta_{N}=1$, let $\bm{\beta}_{n}=(\beta_{n-1},\beta_{n})$ and $\Delta\beta_{n}=\beta_{n}-\beta_{n-1}$. It follows from Lemma˜4 that the sum of the rejection satisfies,
@@
We will first approximate the first term in Equation˜35. Note that since $\Delta W^{\beta}=0$ and $\partial_{\beta^{\prime}}\Delta W^{\bm{\beta}}=\lim_{\beta^{\prime}\to\beta}\Delta W^{\bm{\beta}}/\Delta\beta=:\Delta\dot{W}^{\bm{\beta}}$, we have that $\dot{\alpha}^{\beta}$ satisfies:
@@
For the second term in Equation˜35, we have
@@
By Lemma˜3, $\beta\mapsto\lambda_{\beta}$ is continuously differentiable. Since $\lambda_{\beta}\Delta\beta$ is a right Riemann sum for the integral of $\lambda_{\beta}$ with error:
@@
B.5 Accelerated PT as Vanilla PT on Extended Space
In particular we stress that the distributions $\pi_{\operatorname{ex}}^{n}$ are supported on $\mathcal{X}^{N+1}$ and not $\mathcal{X}$.
Proposition 5.
Proof.
Since the rejection rate only depends on the Radon-Nikodym derivative, it suffices to verify that
which is straightforward from Equation˜36. ∎
B.5.1 Parallelism versus Acceleration Time
The following proposition analyses these local rates for both methods. As in Appendix B.5, the main idea is to find a vanilla PT equivalent for both algorithms. We define
$\tau^{-1}_{\textrm{VA}}(\mu_{0},\ldots,\mu_{K}):=2+2\sum_{k=1}^{K}\frac{r(\mu_{k-1},\mu_{k})}{1-r(\mu_{k-1},\mu_{k})}$
Proposition 6.
Define the sequence of distributions $({\mathbb{S}}_{k})_{k=0}^{K}$ as
Then the following equalities hold
$\tau_{\textrm{TA}}$ $=\tau_{\textrm{VA}}({\mathbb{S}}_{0},{\mathbb{S}}_{K})$ (37)
Proof.
This proposition shows that it is preferable to use the parallel-accelerated method, as the quantity in Equation˜38 is generally greater than that of Equation˜37 thanks to the effect of the bridge between ${\mathbb{S}}_{0}$ and ${\mathbb{S}}_{K}$. However, in practice the time-accelerated method consumes less memory and so might be more suitable in certain circumstances.
Appendix C Schedule Tuning
Under Assumption 1 (efficient local exploration), by Proposition 2, we wish to optimise our schedule $0=\beta_{0}<\ldots<\beta_{N}=1$ to minimise
$\sum_{n=1}^{N}\frac{r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})}{1-r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})},$
in order to maximise the round trip rate of APT. Additionally, by Theorem 2, we have $\sum_{n=1}^{N}r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})\approx\Lambda_{K}$. Therefore, a reasonable proxy objective is to minimise $\sum_{n=1}^{N}r_{n}/(1-r_{n})$ under the constraint that $\sum_{n=1}^{N}r_{n}=\Lambda_{K}$ and $r_{n}>0$. Indeed this is same reasoning employed in (Syed et al., 2022, Section 5) to derive their schedule tuning algorithm.
Thus, following Syed et al. (2022), the solution to the proxy objective is to ensure the $r_{n}$ are constant in $n$. This provides a natural objective to choose our schedule in such a way which results in the theoretical rejection rates $r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})$ being uniform for all $n$. Moreover, while we do not have access to $r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})$, we can estimate these quantities empirically through computing the average rejection rates $\hat{r}_{K}^{n}$ observed in running Algorithm˜1. Specifically, during sampling, we keep track of the average $\hat{\alpha}_{K}^{n}$ of the computed acceptance probabilities $\alpha_{K}^{n}$ for $n=1,\ldots,N$. The estimated rejection rates $\hat{r}_{K}^{n}\approx r(\mathbb{P}_{K}^{n-1},\mathbb{Q}_{K}^{n})$ are then given by $\hat{r}_{K}^{n}=1-\hat{\alpha}_{K}^{n}$.
Finally, to achieve our above goal, we can use the same schedule tuning algorithm from (Syed et al., 2022, Section 5). For completeness, we reproduce this in Algorithm 2.
Algorithm 2 Schedule Tuning
1:Initial schedule $0=\beta_{0}<\ldots<\beta_{N}=1$, number of tuning steps $m$, number of sampling steps $T_{\text{tune}}$;
2:for $i=1,\dots,m$ do
3:  Run Algorithm˜1 for $T_{\text{tune}}$ steps and compute the average rejection rates $\{\hat{r}_{K}^{n}\}_{n=1}^{N}$.
4:  Set $\Lambda_{K}(n)=\sum_{j=1}^{n}r_{K}^{n}$ for $n=1,\ldots,N$ and $\Lambda_{K}(0)=0$.
5:  $S\leftarrow\{(\frac{\Lambda_{K}(n)}{\Lambda_{K}(N)},\beta_{n})\}_{n=0}^{N}$
6:   Fit a monotonically increasing interpolation $\varphi(u)$ (e.g. spline) of the points in $S$ which satisfies the boundary conditions: $\varphi(0)=0$ and $\varphi(1)=1$.
7:  $\beta_{n}\leftarrow\varphi^{-1}\left(\frac{n}{N}\right)$
8:end for
[…]
Appendix D Further Details on Design-Space for Accelerated PT
D.1 Further Details on Flow APT
D.1.1 Work Formula for Deterministic flows